The entrace cover charge for a theme park is $ 1.25 for children and $2.75 for adults. On a particular day the park received 2256 quest paying a total of $ 4089 in cover charge. How kids and adults visited the park?
Answer by ankor@dixie-net.com(12701)

The entrance cover charge for a theme park is $ 1.25 for children and $2.75 for adults. On a particular day the park received 2256 quest paying a total of $ 4089 in cover charge. How kids and adults visited the park?
:
Let a = no. of adults; Let c = no. of children
:
Equation for this statement,"On a particular day the park received 2256 quests"
a + c = 2256
or
c = (2256 - a); use for substitution
:
Equation for this phrase,"paying a total of $4089"
2.75a + 1.25c = 4089
:
Substitute (2256-a) for c in the above equation:
2.75a + 1.25(2256-a) = 4089
:
2.75a + 2820 - 1.25a = 4089
:
2.75a - 1.25a = 4089 - 2820
:
1.5a = 1269
:
a = 1269/1.5
:
a = 846 adults
:
Find children:
c = 2256 - 846
:
c = 1410 children
:
:
Check solution:
2.75(846) + 1.25(1410)
2326.50 + 1762.50 = 4089; our given total